OpenStudy (anonymous):

log 5-log 2x =1 help solving

OpenStudy (anonymous):

\(\log(5)-\log(2x)=1\) like this ?

OpenStudy (anonymous):

I guess. Its written like I wrote it.

OpenStudy (anonymous):

there is a rule: \(\log(A)-\log(B)=\log(A\times B)\). apply this rule to your problem what equation do you get after applying this rule?

OpenStudy (anonymous):

log10x?

OpenStudy (anonymous):

my bad, the rule should say \(\log(A)-\log(B)=\log(A\div B)\)

OpenStudy (anonymous):

log2.5

OpenStudy (anonymous):

\(\displaystyle \log(\frac{5}{2x})\)

OpenStudy (anonymous):

\(\displaystyle \log(\frac{5}{2x})=1\)

OpenStudy (anonymous):

I can't find this i my notes.

OpenStudy (anonymous):

the rule? well, believe me the rule does exist

OpenStudy (anonymous):

I know the rule exist, i can't find a problem like this in my notes. so i'm having difficulty understanding

OpenStudy (anonymous):

it's ok, we can go through it together

OpenStudy (anonymous):

\(\displaystyle \large \log(5)-\log(2x)=1\) this is what we had at first, then we applied the rule: log(A)-log(B)=log(A/B) \(\displaystyle \large \log(\frac{5}{2x})=1\)

OpenStudy (anonymous):

so far so good?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

\(\displaystyle \large \log(\frac{5}{2x})=1\) \(\displaystyle \large \log_{10}(\frac{5}{2x})=1\)

OpenStudy (anonymous):

ths is because unspecific base is always 10

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\(\displaystyle \large \log_{10}(\frac{5}{2x})=1\) -----> \(\displaystyle \large 10^1=\frac{5}{2x}\)

OpenStudy (anonymous):

this is exponential relation with logs

OpenStudy (anonymous):

\(\displaystyle \large 10^1=\frac{5}{2x}\) \(\displaystyle \large 10=\frac{5}{2x}\)

OpenStudy (anonymous):

can you solve fro x from here?

OpenStudy (anonymous):

for* x

OpenStudy (anonymous):

no.

OpenStudy (anonymous):

\(\displaystyle \large 10=\frac{5}{2x}\) \(\displaystyle \large 10\color{blue}{\times \frac{2x}{5}}=\frac{5}{2x}\color{blue}{\times \frac{2x}{5}}\)

OpenStudy (anonymous):

multiply accordingly on both sides

OpenStudy (anonymous):

10x/10x?

OpenStudy (anonymous):

1/4 or 0.25?

OpenStudy (anonymous):

\(\displaystyle \large 10\color{blue}{\times \frac{2x}{5}}=\frac{5}{2x}\color{blue}{\times \frac{2x}{5}}\) \(\displaystyle \large 5 \times 2 \color{blue}{\times \frac{2x}{5}}=\frac{5}{2x}\color{blue}{\times \frac{2x}{5}}\) \(\displaystyle \large \cancel{5} \times 2 \color{blue}{\times \frac{2x}{\cancel{5}}}=\frac{5}{2x}\color{blue}{\times \frac{2x}{5}}\) \(\displaystyle \large \color{blue}{ \frac{4}{5}}x=1\)

OpenStudy (anonymous):

\(\displaystyle \large \color{blue}{ \frac{4}{5}}x=1\) \(\displaystyle \large x=\color{blue}{ \frac{5}{4}}\)

OpenStudy (anonymous):

Thats not an answer?

OpenStudy (anonymous):

X is approximately 0.91970.